# Simultaneous Tri-Modal Medical Image Fusion and Super-Resolution using Conditional Diffusion Model

### v2

@@ Abstract
In clinical practice, tri-modal medical image fusion, compared to the existing dual-modal technique, can provide a more comprehensive view of the lesions, aiding physicians in evaluating the disease’s shape, location, and biological activity. However, due to the limitations of imaging equipment and considerations for patient safety, the quality of medical images is usually limited, leading to sub-optimal fusion performance, and affecting the depth of image analysis by the physician. Thus, there is an urgent need for a technology that can both enhance image resolution and integrate multi-modal information. Although current image processing methods can effectively address image fusion and super-resolution individually, solving both problems synchronously remains extremely challenging. In this paper, we propose TFS-Diff, a simultaneously realize tri-modal medical image fusion and super-resolution model. Specially, TFS-Diff is based on the diffusion model generation of a random iterative denoising process. We also develop a simple objective function and the proposed fusion super-resolution loss, effectively evaluates the uncertainty in the fusion and ensures the stability of the optimization process. And the channel attention module is proposed to effectively integrate key information from different modalities for clinical diagnosis, avoiding information loss caused by multiple image processing. Extensive experiments on public Harvard datasets show that TFS-Diff significantly surpass the existing state-of-the-art methods in both quantitative and visual evaluations. The source code will be available at GitHub.
@@
Multimodal image fusion holds promise for combining information from images of different modalities [14, 3] to obtain a more comprehensive diagnostic view. Currently, image fusion is mainly divided into methods based on deep learning [15, 34] and traditional methods [12, 33]. Deep learning-based methods often use generative adversarial networks (GANs) to simulate the distribution of fused images to obtain high-quality fused images [17]. Although GAN-based methods can generate satisfactory fused images, they suffer from issues such as training instability, mode collapse and lack of interpretability. As an improvement, fusion methods based on Diffusion [20, 7] have been proposed, which generate high-quality images by simulating the diffusion process of restoring images corrupted by noise to clean images, thereby mitigating common problems like training instability and mode collapse in GANs, and their generation process is interpretable. For example, Zhao et al. [35] proposed using DDPM for fusion tasks and employing a hierarchical Bayesian method to model the subproblems of maximum likelihood estimation. However, no deep learning fusion methods for bimodal medical images have emerged, and only a few traditional methods have conducted preliminary research on this problem. For instance, Jie et al. [9] proposed a tri-modal medical image fusion based on an adaptive energy selection scheme and sparse representation, using sparse representation to fuse texture components, and adaptive energy selection scheme to fuse cartoon components. Jie et al.[8] proposed a tri-mode medical image fusion and denoising method based on BitonicX filtering. This method analyzes pixels in terms of gradient, energy, and sharpness to achieve medical image fusion and denoising.
@@ 3.1 Experimental Detail
The dataset covers five different types of registered medical images, including MR-T2/MR-Gad/PET, CT/MR-T2/SPECT, MR-T1/MR-T2/PET, MR-T2/MR-Gad/SPECT, and MR-T1/MR-T2/SPECT. All source images are from the whole brain atlas database of Harvard Medical School[24]. The resolution of the training and testing images is 256x256, which was downsampled using bicubic interpolation to construct super-resolution datasets with different magnification levels (8x, 4x, 2x).The Ground Truth is fused by the BitonicX [8].
@@ 4 Conclusion
4.0.1 Acknowledgement.
This research was supported in part by the National Natural Science Foundation of China under Grant 62201149, in part by the Joint Fund for Basic and Applied Basic Research of Guangdong Province under Grant 2023A1515140077, in part by the Natural Science Foundation of Guangdong Province under Grant 2024A1515011880, in part by the Guangdong Higher Education Innovation and Strengthening of Universities Project under Grant 2023KTSCX127, and in part by the Foshan Key Areas of Scientific and Technological Research Project under Grant 2120001008558, Guangdong, China.

### v3

@@ Abstract
In clinical practice, tri-modal medical image fusion, compared to the existing dual-modal technique, can provide a more comprehensive view of the lesions, aiding physicians in evaluating the disease’s shape, location, and biological activity. However, due to the limitations of imaging equipment and considerations for patient safety, the quality of medical images is usually limited, leading to sub-optimal fusion performance, and affecting the depth of image analysis by the physician. Thus, there is an urgent need for a technology that can both enhance image resolution and integrate multi-modal information. Although current image processing methods can effectively address image fusion and super-resolution individually, solving both problems synchronously remains extremely challenging. In this paper, we propose TFS-Diff, a simultaneously realize tri-modal medical image fusion and super-resolution model. Specially, TFS-Diff is based on the diffusion model generation of a random iterative denoising process. We also develop a simple objective function and the proposed fusion super-resolution loss, effectively evaluates the uncertainty in the fusion and ensures the stability of the optimization process. And the channel attention module is proposed to effectively integrate key information from different modalities for clinical diagnosis, avoiding information loss caused by multiple image processing. Extensive experiments on public Harvard datasets show that TFS-Diff significantly surpass the existing state-of-the-art methods in both quantitative and visual evaluations. Code is available at https://github.com/XylonXu01 /TFS-Diff.
@@
Multimodal image fusion holds promise for combining information from images of different modalities [14, 3] to obtain a more comprehensive diagnostic view. Currently, image fusion is mainly divided into methods based on deep learning [15, 34] and traditional methods [12, 33]. Deep learning-based methods often use generative adversarial networks (GANs) to simulate the distribution of fused images to obtain high-quality fused images [17]. Although GAN-based methods can generate satisfactory fused images, they suffer from issues such as training instability, mode collapse and lack of interpretability. As an improvement, fusion methods based on Diffusion [20, 7] have been proposed, which generate high-quality images by simulating the diffusion process of restoring images corrupted by noise to clean images, thereby mitigating common problems like training instability and mode collapse in GANs, and their generation process is interpretable. For example, Zhao et al. [35] proposed using DDPM for fusion tasks and employing a hierarchical Bayesian method to model the subproblems of maximum likelihood estimation. However, no deep learning fusion methods for tri-modal medical images have emerged, and only a few traditional methods have conducted preliminary research on this problem. For instance, Jie et al. [9] proposed a tri-modal medical image fusion based on an adaptive energy selection scheme and sparse representation, using sparse representation to fuse texture components, and adaptive energy selection scheme to fuse cartoon components. Jie et al.[8] proposed a tri-mode medical image fusion and denoising method based on BitonicX filtering. This method analyzes pixels in terms of gradient, energy, and sharpness to achieve medical image fusion and denoising.
@@ 3.1 Experimental Detail
The dataset covers five different types of registered medical images, including MR-T2/MR-Gad/PET, CT/MR-T2/SPECT, MR-T1/MR-T2/PET, MR-T2/MR-Gad/SPECT, and MR-T1/MR-T2/SPECT. All source images are from the whole brain atlas database of Harvard Medical School[24], We randomly divided the data into 84, 10 and 25 groups as training set, validation set and test set respectively.The resolution of the training and testing images is 256x256, which was downsampled using bicubic interpolation to construct super-resolution datasets with different magnification levels (8x, 4x, 2x).The Ground Truth is fused by the BitonicX [8].
@@ 4 Conclusion
4.0.1 Acknowledgements
This research was supported in part by the National Natural Science Foundation of China under Grant 62201149, in part by the Basic and Applied Basic Research of Guangdong Province under Grant 2023A1515140077, in part by the Natural Science Foundation of Guangdong Province under Grant 2024A1515011880, in part by the Guangdong Higher Education Innovation and Strengthening of Universities Project under Grant 2023KTSCX127, and in part by the Foshan Key Areas of Scientific and Technological Research Project under Grant 2120001008558, China.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.